OpenStudy (anonymous):

Carolyn bought a dozen bagels which included 2 garlic, 3 sesame seed, 1 plain, 4 wheat, and 2 cinnamon sugar. If Carolyn randomly selected a bagel from the bag, what is the probability that she selected a cinnamon sugar bagel?

OpenStudy (anonymous):

2+3+1+4+2=x there are 2 cinnamon sugars, so the answer is 2/x

OpenStudy (anonymous):

no its 1/4 cause you simplify it look the answers are 1/2 1/6 and 1/12

OpenStudy (anonymous):

@danielcb114 aye am i right on this one bro ?

OpenStudy (anonymous):

i totally skipped past the dozen bagel part. so, 2/12=1/6

OpenStudy (anonymous):

lol thats what i did skipped over one of them lol

OpenStudy (anonymous):

ur good im posting a new one to check

OpenStudy (anonymous):

hahah thank you. go for it.

OpenStudy (anonymous):

It's 1/6. There are 12 bagels and 2 cinnamon bagels(my favorite!). So 1/6.